# A Benchmark for Optimal Multi-Modal Multi-Robot Multi-Goal Path Planning with Given Robot Assignment

### v1

@@ Abstract
In many industrial robotics applications, multiple robots are working in a shared workspace to complete a set of tasks as quickly as possible. Such settings can be treated as multi-modal multi-robot multi-goal path planning problems, where each robot has to reach an ordered sequence of goals. Existing approaches to this type of problem solve this using prioritization or assume synchronous completion of tasks, and are thus neither optimal nor complete. We formalize this problem as a single path planning problem and introduce a benchmark encompassing a diverse range of problem instances including scenarios with various robots, planning horizons, and collaborative tasks such as handovers.
Along with the benchmark, we adapt an RRT* and a PRM* planner to serve as a baseline for the planning problems. Both planners work in the composite space of all robots and introduce the required changes to work in our setting.
Unlike existing approaches, our planner and formulation is not restricted to discretized 2D workspaces, supports a changing environment, and works for heterogeneous robot teams over multiple modes with different constraints, and multiple goals.
Videos and code for the benchmark and the planners is available at https://vhartman.github.io/mrmg-planning/.
@@ I Introduction
As adoption of robots increases and simple tasks become more and more automated, it will be more and more important to deploy solutions that are not only able to work longer and cheaper but also are competitive in throughput with humans. In order to achieve this, in many cases, multiple robots need to be used and effectively coordinated in the same workspace: Enabling motion planning for multiple tasks with multiple robots is crucial in order to maximize the usefulness of robots in industrial settings. While workcells with multiple robots exist, the robots typically act independent from each other in order to simplify the programming and avoid dealing with robot-robot interactions.
Most work in continuous multi-robot planning is focusing on single-goal settings where all robots starting from their respective starting points at the same time and reach their respective goals simultaneously [1, 2, 3]. Conversely, in most real use cases, multiple robots need to do multiple tasks in sequence, e.g., welding multiple points in a welding application, or picking and placing multiple things after another in order to sort objects. Since the robots act in the same environment in these scenarios and thus possibly block each other from doing their respective task, we can not formulate the problem at hand as a sequence of path planning problems, but need to solve the multi-robot multi-goal planning problem if we want to find an optimal solution.
The work on multi-modal planning can be seen as multi-goal planning problem. Multi-modal path planning [4, 5, 6, 7] finds paths through sequences of modes, i.e., through variations of a continuous configuration space that occur, e.g., by grasping an object, or moving an object in the workspace. This means that in each mode a different set of constraints is active. Most work on multi-modal path planning for robots only considers single-robot settings.
On the other hand, multi-agent path finding (MAPF) [8, 9] deals with high numbers of agents, but typically considers grid-like 2d environment with disk robots, as e.g., found in warehouses. MAPF was extended to consider multiple goals in the setting of multi agent pickup and delivery (MAPD) [10]. These approaches do not directly transfer to a changing environments and continuous configuration spaces.
@@
  • two probabilistically complete and almost-surely asymptotically optimal planners based on an RRT* planner, and a PRM* planner, respectively.
We open-source an easily accessible multi-modal, multi-robot, multi-goal motion planning benchmark containing 21 different base-scenarios with up to 74 subgoals, that can be further adjusted in difficulty by changing the number of robots and tasks.
Our goal with the planners is not to provide planners with the highest performance possible, but rather to to make the problem accessible and provide baselines.
@@ II Related Work
Most multi robot planning work focusing either on the setting where all robots share the same work and configuration space, or on solving single goal problems. It is commonly assumed that the environment stays completely static during planning, therefore excluding manipulation planning settings. We try to group the different research areas and their respective focus in Table I. In the following, we discuss single-goal multi-robot planning and MAPF more in-depth.
II-A Continuous Multi Robot Planning
A simple approach to multi-robot-planning is planning in the composite space of all robots [14], and applying standard methods such as RRT(*) [19] or PRM(*) [20] in this higher dimensional space. This approach is not scaleable to many robots or high degrees of freedom.
There are various approaches to tackle this limitation. Discrete RRT (dRRT(*) [21, 13]), builds a roadmap for each robot, and combines them to implicitly form a tensor graph. Instead of planning naively on the implicit graph, an RRT-like strategy is proposed to explore this implicit graph. Fast-dRRT [17] applies this approach sequentially on the subproblems to solve multi-goal problems. M* [1] proposes to plan in the separate spaces when possible, and planning jointly only where required, i.e., when conflicts between the single-robot plans arise. Similarly, conflict based search (CBS) [22] plans for each robot individually, and introduces constraints to solve conflicts between single agent plans. CBS originates form the MAPF-setting, but can also be applied to continuous multi robot planning problems [23, 24, 12].
Prioritized planners as proposed in, e.g., [15, 16, 25, 18] achieve more scaleable solvers by not requiring completeness and optimality. Similarly, [26] proposes a scaleable approach that is based on operator splitting and sampling, and while the approach is complete, it is not optimal.
To the best of the authors knowledge, there are no multi-modal multi-robot multi-goal planners that are complete and optimal in the general setting. The prioritized planners can be used to plan for multi goal problems by incrementally generating a path per agent, and considering it as fixed for the later planning problems as done in [18, 16].
@@ II-A Continuous Multi Robot Planning
Compared to the previously introduced works, we are interested in optimally solving multi-robot multi-goal planning problems without any restricting assumptions. Further, we are interested in manipulation planning problems, i.e., problems where the environment changes through the actions of the robot.
II-B Multi Agent Path Finding
Compared to multi robot path planning, multi agent path finding [9] generally refers to planning in grid-like 2D environments with disk-robots, as found in warehouses. A common approach to solve this problem is conflict based [22] or priority based search (PBS) [28]. There are many works that improve upon the initial approach, such as improved conflict based search [29], or bounded suboptimal CBS [30], which scale to 1000s of agents.
Multi agent pickup and delivery (MAPD) refers to the setting where we do not only intend to go from point A to point B, but traveling between multiple points of interest, as part of a single planning problem [10]. While MAPD can be framed as lifelong MAPF problem, i.e., once an agent has finished its task of the MAPF problem, a new goal is assigned to the robot, these solvers are suboptimal [31]. Multi-label A* [11] (MLA*) extends the classic A* algorithm to a sequence of multiple goals, and is able to optimally solve MAPD problems. Both MAPF and MAPD planners typically assume that the environment remains unchanged over the planning duration.
III Multi Robot Multi Goal Path Planning
Informally, our objective is to find a path for each robot that passes through a sequence of goals that minimizes a cost (e.g., minimizing the latest completion time of all robots) while being collision free. Some of the goals might involve multiple robots: a handover of an object between two robots implies a constraint on two robots. A goal could imply a mode transition, i.e., a robot grasping something and thus changing the environment for the remaining planning problem.
@@ III-A Preliminaries
Before formalizing the problem, we introduce the components that are required: the tasks, the concept of a mode, and the state space that we plan in. We follow the work from Thomason [4], and generalize it to multiple robots. Compared to [4], we do not consider task planning.
@@ III-A1 Task
A task $S$ consists of the robots that are assigned to the task, and a set of constraints $g$ [^1] Typically, the constraint $g$ is a single goal pose or a goal region, but can also be a more complex constraint such as a grasp constraint. that need to be fulfilled to consider the task done. A task can have post-conditions that can alter the scene-graph of the environment, i.e., which objects are linked to each other. As example, consider the task ’robot $r_{1}$ grasps object $o_{1}$’: Here, the goal constraints are that the robot is grasping $o_{1}$, and the post-condition is that $o_{1}$ is linked to the end-effector of the robot.
@@ III-A2 Modes
The constraints from a task can be fulfilled by various different poses, which might affect the environment that we plan in. Consider again the grasping-task: the robot is able to fulfill a grasping constraint by grasping from any side. This changes the collision geometry that we need to consider in the rest of the planning problem, and it might influence how the constraints for follow-up tasks can be fulfilled. We use mode to refer to the combination of the discrete task assignment $s$ (which implies a scene-graph), and the poses of all movable objects, defined by the relative transformation to their parent-frames. Compared to related work ([32, 4]), in this work a mode implies what each of the robots is doing.
Figure 2: Examples for the different ways in which a goal sequence can be specified: (left) As a sequence of goals (middle) as a partially ordered sequence and (right) as a per-robot ordered sequence.
Figure 3: Example of a mode-graph that is implied by the dependency graph on the left. The modes with full opacity are an example for how a set of modes can be incrementally be reached. Each mode can be reached with many different parameterizations $\theta_{i}^{\text{k}}$, as indicated for mode [4, 2].
III-A3 Task sequence
A task sequence can be specified either as total ordering, fully determining which tasks come after another, or as dependency graph by specifying which tasks depend on others being fulfilled, which can be processed into a full ordering per robot. We require the terminal task to involve all robots in both the graph and the sequence. Both described specifications can be expressed as dependency graph. We thus describe a task sequence as a dependency graph $G=(S,E)$, with the tasks $S\in\mathcal{S}_{r_{i}}$, and the directed edges between tasks $E\in(\mathcal{S}_{r_{i}}\times\mathcal{S}_{r_{j}})$. Some examples of how a task sequence can be specified are in Fig. 2.
A given (partial) task sequence implies all possible transitions between task assignments (Fig. 3), and thus all valid task assignment sequences. We use $\mathcal{M}$ to denote the set of all valid task assignment sequences that bring us from the start to the goal, and obey the constraints imposed by the dependency graph.
@@ III-A4 State Space
The space in which we describe a path is the composite space of all robots, all objects, and which tasks are currently active per robot:
$\mathcal{Q}=\underbrace{\mathcal{Q}_{r_{1}}\times\cdots\times\mathcal{Q}_{r_{N}}}_{\mathcal{Q}_{R}}\times\mathcal{Q}_{o}\times\underbrace{\mathcal{S}_{r_{1}}\times\cdots\times\mathcal{S}_{r_{N}}}_{\mathcal{S}_{R}}.$ (1)
@@ III-B Problem formulation
Bringing everything together, a multi-modal, multi-robot, multi-goal path planning problem is given by the tuple $(R,\mathcal{Q}_{R},q_{\text{start}},G)$, where $R$ is the set of robots, $\mathcal{Q}_{R}$ is the configuration space of the robots, $q_{\text{start}}$ is the initial state, and $G$ is the dependency graph of the tasks. In the following, we will use $q^{r_{i}}$ for the pose of robot $r_{i}$.
We want to find a collision free path $\pi(t):\mathbb{R}\to\mathcal{Q}_{R}$ and the task assignment sequence $s(t):\mathbb{R}\to\mathcal{S}_{R}$, while minimizing some cost function $c(\cdot)$. The path $\pi$ maps time to the joint robot state $\mathcal{Q}_{R}$, and the mode sequence $s(t)$ maps time to the task assignment $\mathcal{S}_{R}$, which together imply the scene-graph at a time, and thus the poses of all objects.
This can be written as optimization problem:
$\min_{\pi,s}$ $c(\pi,s)$ (2a)
s.t. $\pi(0)=q_{\text{start}}$ (2b)
$g(\pi(t),s(t)\leq 0)\ \ \forall\,t$ (2c)
$s\in\mathcal{M}$ (2d)
$\pi(t)\in\mathcal{Q}_{\text{free}}(s(t))\ \ \forall\,t,$ (2e)
where (2a) is the cost term, (2b) is the initial condition, (2c) and (2d) ensure that the constraints that are imposed by the mode sequence hold, and that the mode sequence is a valid one, and finally (2e) is the constraint for ensuring that the path is collision free.
III-B1 Cost functions
@@
$c(q_{1},q_{2})=(1\!-\!w)\max_{r}||q^{r}_{1}\!-\!q^{r}_{2}||_{2}+w\sum\nolimits_{r}||q^{r}_{1}\!-\!q^{r}_{2}||_{2}.$ (3)
where if $w$ is small, we get a minimum makespan optimization problem (which is ‘regularized’ by the path length), and if $w$ is $1$ we sum the cost of the robots.
@@ IV Planners
We introduce an RRT based planner and a PRM based planner as baselines for the problems we propose. In both these sampling based planners, the space we plan in is the previously defined one. However, instead of directly sampling the (unactuated) object poses, we sample $x=(q,m)$, where $q\in\mathcal{Q}_{R}$ is a pose containing all robots’ poses and $m$ is a mode that determines the task assignment and all object poses. We elaborate on how modes are sampled in Sections IV-A and IV-B. We then proceed in both the PRM* and the RRT* planners as their standard version would, with an adjusted distance function that accounts for the connectivity between the modes. Before discussing the details of the planners, we discuss the distance function used in the planners, and the concept of the informed set [33].
@@ IV-1 Distance functions
$d(x_{1},x_{2})=\begin{cases}d_{\text{mode}}(q_{1},q_{2})&\text{if }m_{1}=m_{2},\\ \infty&\text{otherwise}.\end{cases}$
For the in-mode distance any metric can be chosen; the work in [34] discusses distance metrics in multi-robot problems in more depth. For our planners, we choose the maximum of the per-robot $L_{2}$ distance, as this performed best in our experiments.
IV-2 Locally informed sampling
The informed set [33] can be used to restrict the set of points that we need to consider for planning to the ones that can improve the current solution. The informed set in our problem is
$\mathcal{Q}_{\text{inf}}=\{q\,|\,c_{\text{lb}}(q_{\text{start}},q)+c_{\text{lb}}(q,Q_{\text{goal}})\leq c_{\text{best}}\},$ (4)

### v2

# Sampling-Based Multi-Modal Multi-Robot Multi-Goal Path Planning

@@ Abstract
In many robotics applications, multiple robots are working in a shared workspace to complete a set of tasks as fast as possible. Such settings can be treated as multi-modal multi-robot multi-goal path planning problems, where each robot has to reach a set of goals. Existing approaches to this type of problem solve this using prioritization or assume synchronous task completion, and are thus neither optimal nor complete. We formalize this problem as a single centralized path planning problem and present planners that are probabilistically complete and asymptotically optimal. The planners plan in the composite space of all robots and are modifications of standard sampling-based planners with the required changes to work in our multi-modal, multi-robot, multi-goal setting. We validate the planners on a diverse range of problems including scenarios with various robots, planning horizons, and collaborative tasks such as handovers, and compare the planners against a suboptimal prioritized planner.
Videos and code for the planners and the benchmark is available at https://vhartmann.com/mrmg-planning/.
@@ I Introduction
As adoption of robots increases and simple tasks become more and more automated, it will be increasingly important to deploy solutions that are not only able to work longer and cheaper but are also competitive in throughput with humans. In order to achieve this, in many cases, multiple robots need to be used and effectively coordinated in the same workspace: Enabling motion planning for multiple tasks with multiple robots is crucial in order to maximize the usefulness of robots in industrial settings. While workcells with multiple robots exist, the robots typically act independently from each other in order to simplify the programming and avoid dealing with robot-robot interactions.
Most work in continuous multi-robot planning is focusing on single-goal settings where all robots start moving at the same time and reach their respective goals simultaneously [1, 2, 3]. Conversely, in most real use cases, multiple robots need to do multiple tasks in sequence, e.g., welding multiple points, or picking and placing multiple things after another in order to sort objects. Even when only considering a single pick and place task per robot, each robot needs to reach two goals: The pick, and the place location. Since the robots act in the same environment in these scenarios and thus possibly block each other from doing their tasks, we can not formulate the problem at hand as a sequence of path planning problems, but need to solve the multi-robot multi-goal planning problem if we want to find an optimal solution.
Multi-modal planning can be seen as multi-goal planning problem: Multi-modal path planning [4, 5, 6, 7] finds paths through sequences of modes, i.e., through variations of a continuous configuration space that occur, e.g., by grasping an object, or moving an object in the workspace. This means that in each mode a different set of constraints is active. Most work on multi-modal path planning for robots only considers single-robot settings.
On the other hand, multi-agent path finding (MAPF) [8, 9] deals with high numbers of agents, but typically considers grid-like 2d environments with a homogeneous group of disk robots, as found, e.g., in warehouses. MAPF was extended to consider multiple goals in the multi-agent pickup and delivery (MAPD) setting [10]. These approaches do not transfer easily to changing environments and continuous configuration spaces.
@@
  • adaptions of probabilistically complete and almost-surely asymptotically optimal planners to this setting.
We compare the planners to a prioritized planner on a multi-modal, multi-robot, multi-goal motion planning benchmark containing 21 different base-scenarios with up to 74 subgoals, that can be further adjusted in difficulty by changing the number of robots and tasks, which we open-source.
@@ II Related Work
Most multi-robot planning work is focusing either on the setting where all robots share the same work and configuration space, or on solving single goal problems. It is commonly assumed that the environment stays completely static during planning, therefore excluding manipulation planning settings. We group the different research areas and their respective focus in Table I. In the following, we discuss single-goal multi-robot planning and MAPF more in-depth.
II-A Continuous Multi-Robot Planning
A simple approach to multi-robot planning is planning in the composite space of all robots [14], and applying standard methods such as RRT(*) [19] or PRM(*) [20] in this higher dimensional space. This approach is not scalable to many robots or high degrees of freedom.
There are various approaches to tackle this limitation. Discrete RRT (dRRT(*) [21, 13]), builds a roadmap for each robot, and combines them to implicitly form a tensor graph. Instead of planning naively on the implicit graph, an RRT-like strategy is proposed to explore this implicit graph. Fast-dRRT [17] applies this approach sequentially on the subproblems to solve multi-goal problems. M* [1] plans in the separate spaces when possible, and plans jointly only where required, i.e., when conflicts between the single-robot plans arise. Similarly, conflict based search (CBS) [22] plans for each robot individually, and introduces constraints to solve conflicts between single agent plans. CBS originates form the MAPF-setting, but can also be applied to continuous multi-robot planning problems [23, 24, 12].
Prioritized planners as proposed in, e.g., [15, 16, 25, 18] achieve more scalable solvers by not requiring completeness and optimality. Similarly, [26] proposes a scalable approach that is based on operator splitting and sampling, and while the approach is complete, it is not optimal.
The prioritized planners can be used to plan for multi-goal problems by incrementally generating a path per agent, and considering it as fixed for the later planning problems as done in [18, 16]. To the best of the authors’ knowledge, there are no multi-modal multi-robot multi-goal planners that are complete and optimal in the general setting.
@@ II-A Continuous Multi-Robot Planning
Compared to the discussed works, we are interested in optimally solving multi-robot multi-goal planning problems without any restricting assumptions. Further, we are interested in manipulation, i.e., problems where the environment changes through the actions of the robot.
II-B Multi-Agent Path Finding
MAPF [9] generally refers to planning in grid-like 2D environments with disk-robots, as found in warehouses. A common approach to solve this problem is conflict based [22] or priority based search (PBS) [28]. There are many works that improve upon the initial approach, such as improved CBS [29], or bounded suboptimal CBS [30], which scale to 1000s of agents.
MAPD refers to the setting where we do not only intend to go from point A to point B, but traveling between multiple points of interest, as part of a single planning problem [10, 31]. While MAPD can be framed as lifelong MAPF problem, i.e., once an agent has finished its task of the MAPF problem, a new goal is assigned to the robot, these solvers are suboptimal [32]. Multi-label A* [11] (MLA*) extends the classic A* algorithm to a sequence of multiple goals, and is able to optimally solve MAPD problems. Both MAPF and MAPD planners typically assume that the environment remains unchanged over the planning duration.
III Multi-Robot Multi-Goal Path Planning
Informally, our objective is to find a collision free path for each robot that passes through a sequence of goals that minimizes a cost (e.g., the latest completion time of all robots). Goals might involve multiple robots: A handover of an object between two robots implies a constraint on two robots. A goal could additionally imply a mode transition, i.e., a robot grasping something and thus changing the environment for the remaining planning problem.
@@ III-A Preliminaries
Before formalizing the problem, we introduce the components that are required: The tasks, the concept of a mode, and the state space that we plan in. We follow the work from Thomason et al. [4], and generalize it to multiple robots. Compared to [4], task sequencing and allocation happens implicitly in the planning.
@@ III-A1 Task
A task $S$ consists of the robots that are assigned to the task, and a set of constraints gg [^1] Typically, the constraint $g$ is a single goal pose or a goal region, but can also be a more complex constraint such as a grasp constraint. that need to be fulfilled to consider the task done. A task can have post-conditions that can alter the scene-graph of the environment, i.e., which objects are linked to each other. As example, consider the task ’robot $r_{1}$ grasps object $o_{1}$’: Here, the goal constraints are that the robot is grasping $o_{1}$, and the post-condition is that $o_{1}$ is linked to the end-effector of the robot.
@@ III-A2 Modes
The constraints of a task can be fulfilled by a set of poses, where the chosen pose might affect the environment that we plan in. Consider again the grasping-task: The robot is able to fulfill a grasping constraint by grasping from any side, which changes the collision geometry that we need to consider in the rest of the planning problem and it might influence how the constraints for follow-up tasks can be fulfilled. We use mode $m\in\mathcal{M}=\mathcal{S}\times\mathcal{Q}_{o}$ to refer to the combination of the discrete task assignment $s$ (which implies a scene-graph), and the poses of all movable objects $q_{o}\in\mathcal{Q}_{o}$, defined by the relative transformation to their parent-frames.
Figure 2: Examples for the different ways in which a task ordering can be specified: (left) As a sequence of tasks (middle) as a partially ordered sequence and (right) as a set of unassigned and unordered tasks (i.e., the order of task groups 1 and 2 is exchangeable).
Figure 3: Example of a mode-graph that is implied by the ordering on the left. Modes with full opacity show an example of how a set of modes can be incrementally reached. Each mode can be reached with different parameterizations $\theta_{i}^{\text{k}}$ (e.g., different grasps), as indicated for mode (4, 2).
III-A3 Task Order Specification
A task sequence is induced by an oracle $\mathcal{O}(m):\mathcal{M}\rightarrow\mathcal{S}^{*}$ which maps the current assignment and state of the environment to the list of possible task assignments $\mathcal{S}^{*}$. In practice, we describe the ordering of tasks as (partially connected) dependency graph $G=(S,E)$, with the tasks $S\in\mathcal{S}_{r_{i}}$, and the directed edges between tasks $E\in(\mathcal{S}_{r_{i}}\times\mathcal{S}_{r_{j}})$. We require the terminal task to involve all robots. Some examples of how a task sequence can be specified are illustrated in Fig. 2. A (partial) task sequence implies all possible transitions between task assignments (Fig. 3), and thus all logically valid task assignment sequences. We use $\mathcal{T}$ to denote the set of all valid task assignment sequences that bring us from the start to the goal and obey the constraints imposed by the oracle.
@@ III-A4 State Space
We describe a path in the composite space of all robots, all objects, and which tasks are currently active per robot:
$\mathcal{Q}=\underbrace{\mathcal{Q}_{r_{1}}\times\cdots\times\mathcal{Q}_{r_{N}}}_{\mathcal{Q}_{R}}\times\mathcal{Q}_{o}\times\underbrace{\mathcal{S}_{r_{1}}\times\cdots\times\mathcal{S}_{r_{N}}}_{\mathcal{S}_{R}}.$
@@ III-B Problem formulation
Bringing everything together, a multi-modal, multi-robot, multi-goal path planning problem is given by the tuple $(R,\mathcal{Q}_{R},q_{\text{start}},m_{\text{start}},\mathcal{O})$, where $R$ is the set of robots, $\mathcal{Q}_{R}$ is the configuration space of the robots, $q_{\text{start}}$ is the initial state, $m_{\text{start}}$ is the start mode, and $\mathcal{O}$ is the oracle giving us the possible next task assignments from the current mode. In the following, we will use $q^{r_{i}}$ for the pose of robot $r_{i}$.
We want to find a collision free path $\pi(t):\mathbb{R}\to\mathcal{Q}_{R}$ and the task assignment sequence $s(t):\mathbb{R}\to\mathcal{S}_{R}$, that minimizes the cost function $c(\cdot)$. The path $\pi$ maps time to the composite robot state $\mathcal{Q}_{R}$, and the mode sequence $s(t)$ maps time to the task assignment $\mathcal{S}_{R}$, which together imply the scene-graph at a time, and thus the poses of all objects.
Cost functions
@@
[…]
where if $w$ is small, we get a minimum makespan optimization problem (which is ‘regularized’ by the path length), and if $w$ is $1$ the total cost is the sum of all robot path lengths.
@@ IV Planners
We now introduce the necessary adaptions to four sampling-based planners (RRT*, PRM* [33], AIT* and EIT* [34]). We refer the reader to the original works for a detailed explanation of how the algorithms work. In all these planners, the space we plan in is the previously defined one. However, instead of directly sampling the (unactuated) object poses, we sample $x=(q,m)$, where $q\in\mathcal{Q}_{R}$ is a pose containing all robots’ poses and $m$ is a mode that determines the task assignment and all object poses. Generally, the planners then proceed as their standard version would, with an adjusted distance function that accounts for the connectivity between the modes. Before discussing any specific adjustments required in the planners in Section IV-A, we discuss the distance function used in the planners, and the concept of the informed set [35] as well as mode sampling.
@@ IV-1 Distance functions
$d(x_{1},x_{2})=\begin{cases}d_{\text{mode}}(q_{1},q_{2})&\text{if }m_{1}=m_{2},\\ 0&\text{if }q_{1}=q_{2}\text{ and }m_{2}\in\text{succ}(m_{1})\\ \infty&\text{otherwise},\end{cases}$
where we use $\text{succ}(m)$ to indicate the set of successor modes of mode $m$. For the in-mode distance any metric can be chosen; the work in [36] discusses distance metrics in multi-robot problems in more depth. For our planners, we choose the maximum of the per-robot $L_{2}$ distance, as this performed best in our experiments.
IV-2 Mode sampling
As previously introduced, we sample configurations $(q,m)$ from the combined configuration and mode space. In all planners, we store the list of modes $M$ that we reached so far, and expand it whenever we reach a new mode (i.e., whenever we reach a configuration that fulfills the constraints $g$ of a task). Instead of uniformly sampling from the set $M$, we maintain the most recently added mode $F_{-1}$ and the set of modes $F$ that we did not yet expand (the frontier). We introduce a mode sampling strategy specifically tailored to our multi-goal multi-modal problem:
$m\sim\begin{cases}\mathcal{U}(F)&\text{with probability }p(1-\epsilon)\\ \mathcal{U}(F_{-1})&\text{with probability }p\epsilon\\ \mathcal{U}(M\setminus F)&\text{with probability }1-p\end{cases}$
where $p\in[0,1]$ and $\epsilon\in[0,1]$ are weighting parameters, and $\mathcal{U(\cdot)}$ denotes a uniform distribution over a finite set. If $p=1$ and $\epsilon=1$, we reach a fully greedy version, resulting in the most direct progression possible towards the terminal mode.
We refer to the frontier sampling strategy if $p>0.5$ holds. This biases the planner towards the unexpanded modes (i.e., frontiers) compared to the fully greedy approach, while still maintaining a greedy progression toward the terminal state [^2] A fully greedy strategy is not probabilistically complete in general.. For $p=0.5$ and $\epsilon=0$, $M$ is sampled uniformly.
IV-3 Locally informed sampling
Once we found a solution to our planning problem, the informed set [35] can be used to restrict the points that we need to consider for planning to the points that can improve the current solution. The informed set in our problem is
$\mathcal{Q}_{\text{inf}}=\{q\,|\,c_{\text{lb}}(q_{\text{start}},q)+c_{\text{lb}}(q,Q_{\text{goal}})\leq c_{\text{best}}\},$